A piece of wire is shaped to enclose an equilateral triangle whose area is 27.71cm^2. It is then reshaped to enclose a rectangle whose length is 9cm., find the area of the rectangle
Found 2 solutions by rothauserc, Theo:
Answer by rothauserc(4718)

let s be the length of a side of the equilateral triangle
:
area of equilateral triangle in terms of s is
:
Area = s^2 * square root(3)/4
:
Note this formula comes from the fact that the base is s/2 and height = s * square root(3)/2 for an equilateral triangle
:
27.71 = s^2 * square root(3)/4
:
s^2 = (4 * 27.27)/square root(3) = 62.9774
:
s = square root(62.9774) is approximately 7.94 cm
:
Perimeter of the equilateral triangle is 3 * 7.94 = 23.82 cm
:
Perimeter of rectangle is 2l + 2w where l is length and w is width
:
we are given that l=9
:
2(9) +2w = 23.82
:
2w = 23.82 -18 = 5.82
:
w = 2.91 cm
:
Area of rectangle(A) = l * w
:
*************************
A = 9 * 2.91 = 26.19 cm^2
*************************
:

Answer by Theo(13342)

the area of the equilateral triangle is 27.71 square centimeters.

the formula for the area of an equilateral triangle is A = sqrt(3) / 4 * S^2, where S is the side of the equilateral triangle.

this formula becomes 27.71 = sqrt(3) / 4 * S^2

solve for S to get S = sqrt(27.71 * 4 / sqrt(3)) = 7.99959398.

that's the length of one side of the triangle.

the perimeter of the triangle is 3 times that = 23.99878194.

that's the length of the wire that was used to shape the triangle.

the length of the rectangle is 9.

the formula for the perimeter of the rectangle is 2 * (L + W) = 23.99878194.

divide both sides of that equation by 2 to get L + W = 23.99878194 / 2 = 11.99939097.

since L = 9, W must be equal to 11.99939097 minus 9 = 2.999390969.

you have:

L = 9
W = 2.999390969

the area of the rectangle is equal to L * W = 9 * 2.999390969 = 26.99451872 square centimeters.